# Constrained least squares simplicial-simplicial regression

### v2

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. For this setting, constrained least squares, where the regression coefficients themselves lie within the simplex, is proposed. The model is transformation-free but the adoption of a power transformation is straightforward, it can treat more than one compositional datasets as predictors and offers the possibility of weights among the simplicial predictors. Among the model’s advantages are its ability to treat zeros in a natural way and a highly computationally efficient algorithm to estimate its coefficients. Resampling based hypothesis testing procedures are employed regarding inference, such as linear independence, and equality of the regression coefficients to some pre-specified values. The performance of the proposed technique and its comparison to an existing methodology that is of the same spirit takes place using simulation studies and real data examples.
@@ 1 Introduction
Tsagris et al. (2011) categorized the compositional data analysis approaches into two main categories, the raw data approach and the log-ratio approach. A perhaps better classification would be the raw data and the transformation-based approaches. Moving along the lines of the raw data approach the paper proposes the use of the same transformation-free linear regression model, as in Fiksel et al. (2022), when both the response and the predictor variables are simplicial. However, the adoption of a power transformation in the simplicial response generalizes the model. The regression coefficients are estimated via simplicial constrained least squares (SCLS) and as the name implies, least squares is the loss function used to estimate the regression coefficients which are constrained to lie on the simplex. This in turn implies that the expected value of the simplicial response can be expressed as a Markov transition from the simplicial predictor. The proposed SCLS model allows for more than one simplicial predictor, further allows the possibility of assigning weights to the simplicial predictors, and treats zero values naturally, in both the simplicial response and the predictor variables. The assumption of linear independence between the simplicial variables, and hypotheses regarding the mateix of regression coefficients can be tested using resampling techniques. Evidently, the SCLS is similar in spirit to the TFLR, but they have different loss (or objective functions). The TFLR model employs the Expectation-Maximization (EM) algorithm, whereas the SCLS model is based on quadratic programming, thus it enjoys a really low computational cost.
@@
The rest of the paper is structured as follows. Section 2 reviews some simplicial-simplicial regression models, while Section 3 introduces the SCLS model, and discusses several cases related to the TFLR model as well. Section 4 contains Monte-Carlo simulation studies comparing the SCLS to the TFLR model in terms of a) type I and type II errors of the linear independence assumption, b) discrepancy of the estimated regression coefficients, and c) computational cost. The SCLS model is then applied to real data for illustration, and comparison to the TFLR model, purposes, while the last section concludes the paper.
@@ 2.1 Log-ratio simplicial-simplicial regression models
where ${\bf v}=\left(v_{1},\ldots,v_{D-1}\right)\in\mathbb{R}^{D-1}$. Note that the common divisor, $u_{1}$, need not be the first component and was simply chosen for convenience.
@@
where $D_{r}$ and $D_{p}$ denote the number of components of the response and compositional variables, respectively.
@@ 2.2 Kullback-Leibler divergence principal component regression
While the power transformed vector ${\bf w}_{\alpha}$ in Eq. (5) remains in the simplex $\mathbb{S}^{D-1}$, ${\bf z}_{\alpha}$ in Eq. (6) is mapped onto a subset of $\mathbb{R}^{D-1}$. Furthermore, as $\alpha\rightarrow 0$, Eq. (6) converges to the ilr transformation [^3] The scaling factor $D$ exists to assist in the convergence. in Eq. (4) (Tsagris et al., 2016), provided no zero values exist in the data. For convenience purposes, $\alpha$ is generally taken to be between $-1$ and $1$, but when zeros occur in the data, $\alpha$ must be restricted to be strictly positive.
@@ 2.3 The TFLR model
The TFLR model (Fiksel et al., 2022) relates the simplicial response to the simplicial predictor via a linear transformation
@@ 3.1 The SCLS model and quadratic programming
The SL(B) though (13) is minimized via quadratic matrix programming, but it can be formulated under the quadratic programming framework [^6] The SCLS model along with all relevant functions used throught this paper has been implemented in the R package Compositional (Tsagris et al., 2024) which makes use of the R package quadprog (Turlach et al., 2019) that has implemented the algorithm of Goldfarb and Idnani (1983) to minimize Eq. (13)., as in the constrained minimization formula in Eq. (14). The $\bf D$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the ${\bf X}^{\top}{\bf X}$ in the following manner
@@
The ${\bf A}^{\top}$ matrix can be broken down two three sub-matrices, the $D_{p}\times D_{rp}$ ${\bf A}_{1}$, the $D_{rp}\times D_{rp}$ ${\bf A}_{2}$ and the the $D_{rp}\times D_{rp}$ ${\bf A}_{3}$, one stacked under the other,
${\bf A}^{\top}=\left(\begin{array}[]{c}\Huge{\bf A}_{1}^{\top}\\ {\bf A}_{2}^{\top}\\ {\bf A}_{3}^{\top}\end{array}\right)={\bf A}^{\top}=\left(\begin{array}[]{ccc}{\bf I}_{D_{r}}&\ldots&{\bf I}_{D_{r}}\\ &{\mbox{$\textbf{I}_{D_{rp}}$}}&\\ &{\mbox{$-\textbf{I}_{D_{rp}}$}}&\end{array}\right).$
@@
The $\bm{b}$ relates to the $\bf B$ matrix via the vectorization operation $\bm{b}=\mathop{\operator@font vec}\nolimits\left({\bf B}\right)$, that is, each column of the matrix is stacked one under the other. Finally, the $\bm{b}_{0}$ vector contains the $D_{p}$-dimensional vector of 1s, $\bm{j}_{D_{p}}$, that corresponds to the unity sum constraint of the rows of ${\bf B}$ (linked to the ${\bf A}_{1}^{\top}$ matrix), the $D_{rp}$-dimensional vector of 0s, $\bm{0}_{D_{rp}}$, corresponding to the fact that all elements $B_{jk}$ of ${\bf B}$ are non-negative (linked to the ${\bf A}_{2}^{\top}$ matrix) and finally the $D_{rp}$-dimensional vector of -1s, $-\bm{j}_{D_{rp}}$, corresponding to the fact that all elements $B_{jk}$ of ${\bf B}$ are at bounded by unity from above (linked to the ${\bf A}_{3}^{\top}$ matrix)
@@ 3.2 Interpretation of the regression coefficients
Interpretation of the resulting estimated coefficients is a crucial aspect of a regression model if one is interested in making inference about them [^7] Hypothesis testing for the parameters relies upon resampling techniques (presented next) due to the lack of parametric assumptions imposed on the coefficients. . Tsagris et al. (2023) for instance proposed some non-parametric regression models that do not estimate coefficients, and thus visualized the (non-linear) effects of the predictor variables graphically. SCLS on the other hand yields regression coefficients and being a transformation free regression model, interpretation of its regression coefficients has similarities to the interpretation of the coefficients of the regression model with the alr (2) transformation. In the latter, a change in a predictor variable refers to a relative change in the components of the simplicial response, and in the SCLS, the interpretation is somewhat similar. If $x_{j}$ increases by $\delta$, while $x_{l}$ decreases by the same amount, ceteris paribus, the expected change of $\bf y$ is equal to $\delta\left({\bf B}_{j}-{\bf B}_{l}\right)$, where $j$ and $l$ denote rows. Assume for instance the following form of the matrix of regression coefficients
@@ 3.3 Visualization of the regression coefficients
In the case of a 3-part simplicial response, the coefficients can also be visualized using the ternary diagram, as in Figure 1, that contains the 4 rows of the $\bf B$ matrix in Eq. (25). The interpretation of the ternary diagram is as follows. A point close to a vertex indicates high proportion in that vertex-associated component, while a point close to an edge opposite a vertex, indicates a low value in the vertex-associated component and finally a point that lies close to the barycentre of the triangle indicates almost equal values in all components. The third value of the second row of the matrix (${B}_{23}$) is equal to 0.60 and thus closer to the top vertex, whereas the third row values are similar, and hence B3 lies close to the barycentre of the triangle. When the (row-)coefficients of a component of the simplicial predictor take values to the barycenter, i.e. $\left(1/D_{r},\ldots,1/D_{r}\right)$ this indicates that the particular component of the simplicial predictor does not carry too much statistical information. On the contrary, the further the coefficients associated with a component of the simplicial predictor are from the barycentre, the more information this component carries. In this simple example, the third component does not seem very important, because the third row coefficients (B3) are close to the barycentre. The second component (B2) of the simplicial predictor on the other hand, is probably the most important component, because its vector of coefficients lie the furthest from the barycentre among the other three vectors of coefficients.
@@ 3.5 A test for specific values imposed on regression coefficients
A sub-composition stems from a composition by removing at least one component and rescaling the rest of the values to some to unity. The resulting null hypothesis in this regression setting is $B_{j1}=\ldots=B_{jD_{r}}=0$ for some $j$, $1\leq j\leq D_{p}$. However, there is an associated problem with this, the rescaling part, which in order to be valid the property of sub-compositional coherence [^9] This property refers to the invariance of the results when the full composition or the sub-composition is used. (Aitchison, 2003) should be met and in this case it is not. However, the theoretical implications of the sub-compositionality are not known, and hence this test perhaps requires more investigation.
@@ 3.8.1 AR(1) model
In the AR(1) model formulation, the current expected simplicial response is a linear function of its lagged time response, $E({\bf Y}_{t}|{\bf Y}_{t-1})={\bf Y}_{t-1}{\bf B}$, thus the matrix $\bf B$ can be interpreted as the matrix of transition probabilities of the states of the simplicial response at time $t-1$. The linear independence test in this case translates to testing the assumption of no auto-correlation. The case of allowing for more than one simplicial predictors extends the 1 lag to p lags, thus extend the AR(1) model to the AR(p) model. However, there is no theory to this end, what are implications of such a model (i.e. unit root testing, etc.) and thus this requires more research.
@@ 3.8.2 Categorical predictor or response
The case of a categorical predictor can be addressed via the SCLS. In that case, each row of the simplicial predictor is written as ${\bf x}_{i}={\bf e}_{j}$, where ${\bf e}_{j}$ denotes a vector with 0 $D_{p}$ elements except its $j$-th element that is 1101010This could also be seen as an extreme case of compositional data, where one component takes a value of 1 and all other component have zero values., yielding a one-way analysis of variance type of model.
The linear independence test translates to to testing the assumption of equal simplicial population means. The latter case was studied by Tsagris et al. (2017) using parametric and non-parametric procedures. However, simulation studies (not presented) replicating a case scenario from Tsagris et al. (2017) comparing the SCLS and TFLR with the bootstrap version of the two sample James test showed that the latter is size correct, but the former two produced inflated type I errors. Hence, in contrast to the suggestion made by Fiksel et al. (2022), there is evidence against the use of SCLS and TFLR models for the purpose of hypothesis testing of equality of simplicial means.
@@
3.9 Alternative implementation of the TFLR model and generalizations thereof
@@
The simplicial response can be vectorized ${\bf y}^{*}=\mathop{\operator@font vec}\nolimits({\bf Y})$ and then estimate the coefficients of a logistic regression model of ${\bf y}^{*}$ on ${\bf X}^{*}$ with the identity as the link function, no constant term and imposing the same constraints as in the SCLS. This strategy is the same as the in one used in the seemingly unrelated regression equations (SURE) model (Zellner, 1962). The SURE idea can also be used if one wants to substitute the square loss (13), or the KLD loss (12) with the $L_{1}$ norm loss and thus perform a simplicial constrained multivariate median regression (or spatial median regression (Chakraborty, 2003)). The median regression has been implemented in the quantreg R package (Koenker, 2023) and that implementation allows for inequality constraints and hence the constraints will be nearly the same to those used in Eq. (14).
@@ 4.2 Type II error
The estimated powers for both models are presented in Table 2. For the small sample sized data, TFLR seems to produce higher estimated power levels than the SCLS, but as soon as the sample size increases their estimated power is equal.

### v3

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. For this setting, constrained least squares, where the regression coefficients themselves lie within the simplex, is proposed. The model is transformation-free but the adoption of a power transformation is straightforward, it can treat more than one compositional datasets as predictors and offers the possibility of weights among the simplicial predictors. Among the model’s advantages are its ability to treat zeros in a natural way and a highly computationally efficient algorithm to estimate its coefficients. Resampling based hypothesis testing procedures are employed regarding inference, such as linear independence, and equality of the regression coefficients to some pre-specified values. The strategy behind the formulation of the new model is implemented is related to an existing methodology, that is of the same spirit, showcasing how other similar models can be employed as well. Finally, the performance of the proposed technique and its comparison to the existing methodology takes place using simulation studies and real data examples.
@@ 1 Introduction
Tsagris et al. (2011) categorized the compositional data analysis approaches into two main categories, the raw data approach and the log-ratio approach. A perhaps better classification would be the raw data and the transformation-based approaches. Moving along the lines of the raw data approach the paper proposes the use of the same transformation-free linear regression model, as in Fiksel et al. (2022), when both the response and the predictor variables are simplicial. However, the adoption of a power transformation in the simplicial response generalizes the model. The regression coefficients are estimated via simplicial constrained least squares (SCLS) and as the name implies, least squares is the loss function used to estimate the regression coefficients which are constrained to lie on the simplex. This in turn implies that the expected value of the simplicial response can be expressed as a Markov transition from the simplicial predictor. The proposed SCLS model allows for more than one simplicial predictor, further allows the possibility of assigning weights to the simplicial predictors, and treats zero values naturally, in both the simplicial response and the predictor variables. The assumption of linear independence between the simplicial variables, and hypotheses regarding the matrix of regression coefficients can be tested using resampling techniques. Evidently, the SCLS is similar in spirit to the TFLR, but they have different loss (or objective functions). Lastly, the TFLR model employs the Expectation-Maximization (EM) algorithm, whereas the SCLS model is based on quadratic programming, thus it enjoys a really low computational cost.
@@
The rest of the paper is structured as follows. Section 2 reviews some simplicial-simplicial regression models, while Section 3 introduces the SCLS model, and discusses several cases related to the TFLR model as well. Section 4 contains Monte-Carlo simulation studies comparing the SCLS to the TFLR model in terms of a) type I and type II errors of the linear independence assumption, b) discrepancy of the estimated regression coefficients, and c) computational cost. The SCLS model is then applied to real data for illustration, and comparison to the TFLR model, purposes (Section 5), while the last section concludes the paper.
@@ 2.1 Log-ratio simplicial-simplicial regression models
where ${\bf v}=\left(v_{1},\ldots,v_{D-1}\right)\in\mathbb{R}^{D-1}$. Note that the common divisor, $y_{1}$, need not be the first component and was simply chosen for convenience.
@@
where $D_{r}$ and $D_{p}$ denote the number of components of the response and compositional variables, respectively, and ${\bf v}_{k}$ and ${\bf z}_{0k}$ denote the $k$-variable of the alr (2) and ilr (4) transformed simplicial response, respectively.
@@ 2.2 Kullback-Leibler divergence principal component regression
While the power transformed vector ${\bf w}_{\alpha}$ in Eq. (5) remains in the simplex $\mathbb{S}^{D-1}$, ${\bf z}_{\alpha}$ in Eq. (6) is mapped onto a subset of $\mathbb{R}^{D-1}$. Furthermore, as $\alpha\rightarrow 0$, Eq. (6) converges to the ilr transformation [^3] The scaling factor $D$ exists to assist in the convergence. See Tsagris et al. (2016) for more information. in Eq. (4) (Tsagris et al., 2016), provided no zero values exist in the data. For convenience purposes, $\alpha$ is generally taken to be between $-1$ and $1$, but when zeros occur in the data, $\alpha$ must be restricted to be strictly positive.
@@ 2.3 The TFLR model
The TFLR model (Fiksel et al., 2022) relates the $k$-th simplicial response vector to the simplicial predictor via a linear transformation
@@ 3.1 The SCLS model and quadratic programming
The SL(B) though (13) is minimized via quadratic matrix programming, but it can be formulated under the quadratic programming framework [^6] The SCLS model, along with all relevant functions used throughout this paper, has been implemented in the R package Compositional (Tsagris et al., 2024) which makes use of the R package quadprog (Turlach et al., 2019) that has implemented the algorithm of Goldfarb and Idnani (1983) to minimize Eq. (13). Note that this implementation allows for equality constrains as well., as in the constrained minimization formula in Eq. (14). The $\bf D$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the ${\bf X}^{\top}{\bf X}$ in the following manner
@@
The ${\bf A}^{\top}$ matrix can be broken down two three sub-matrices, the $D_{rp}\times D_{p}$ ${\bf A}_{1}$, the $D_{rp}\times D_{rp}$ ${\bf A}_{2}$ and the the $D_{rp}\times D_{rp}$ ${\bf A}_{3}$, one stacked under the other,
${\bf A}^{\top}=\left(\begin{array}[]{c}\Huge{\bf A}_{1}^{\top}\\ {\bf A}_{2}^{\top}\\ {\bf A}_{3}^{\top}\end{array}\right)={\bf A}^{\top}=\left(\begin{array}[]{ccc}{\bf I}_{D_{p}}&\ldots&{\bf I}_{D_{p}}\\ &{\mbox{$\textbf{I}_{D_{rp}}$}}&\\ &{\mbox{$-\textbf{I}_{D_{rp}}$}}&\end{array}\right).$
@@
The $\bm{b}$ relates to the $\bf B$ matrix via the vectorization operation $\bm{b}=\mathop{\operator@font vec}\nolimits\left({\bf B}\right)$ and the $\bf d$ vector in Eq. (14) is the vectorized matrix ${\bf X}^{\top}{\bf Y}$, ${\bf d}={\bf X}^{\top}{\bf Y}$. Finally, the $\bm{b}_{0}$ vector contains the $D_{p}$-dimensional vector of 1s, $\bm{j}_{D_{p}}$, that corresponds to the unity sum constraint of the rows of ${\bf B}$ (linked to the ${\bf A}_{1}^{\top}$ matrix), the $D_{rp}$-dimensional vector of 0s, $\bm{0}_{D_{rp}}$, corresponding to the fact that all elements $B_{jk}$ of ${\bf B}$ are non-negative (linked to the ${\bf A}_{2}^{\top}$ matrix) and finally the $D_{rp}$-dimensional vector of -1s, $-\bm{j}_{D_{rp}}$, corresponding to the fact that all elements $B_{jk}$ of ${\bf B}$ are at bounded by unity from above (linked to the ${\bf A}_{3}^{\top}$ matrix)
@@ 3.2 Interpretation of the regression coefficients
Interpretation of the resulting estimated coefficients is a crucial aspect of a regression model if one is interested in making inference about them [^7] Hypothesis testing for the parameters relies upon resampling techniques (presented next) due to the lack of parametric assumptions imposed on the coefficients. . Tsagris et al. (2023) for instance, proposed some non-parametric regression models that do not estimate coefficients, and thus visualized the (non-linear) effects of the predictor variables graphically. SCLS on the other hand yields regression coefficients and being a transformation free regression model, interpretation of its regression coefficients has similarities to the interpretation of the coefficients of the regression model with the alr (2) transformation. In the latter, a change in a predictor variable refers to a relative change in the components of the simplicial response, and in the SCLS, the interpretation is somewhat similar. If $x_{j}$ increases by $\delta$, while $x_{l}$ decreases by the same amount, ceteris paribus, the expected change of $\bf y$ is equal to $\delta\left({\bf B}_{j}-{\bf B}_{l}\right)$, where $j$ and $l$ denote rows. Assume for instance the following form of the matrix of regression coefficients
@@ 3.3 Visualization of the regression coefficients
In the case of a 3-part simplicial response, the coefficients can also be visualized using the ternary diagram, as in Figure 1, that contains the 4 rows of the $\bf B$ matrix in Eq. (25). The interpretation of the ternary diagram is as follows. A point close to a vertex indicates high proportion in that vertex-associated component, while a point close to an edge opposite a vertex, indicates a low value in the vertex-associated component and finally a point that lies close to the barycentre of the triangle indicates almost equal values in all components. The third value of the second row of the matrix (${B}_{23}$) is equal to 0.60 and thus closer to the top vertex, whereas the third row values are similar, and hence B3 lies close to the barycentre of the triangle. When the (row-)coefficients of a component of the simplicial predictor take values equal to the barycenter, i.e. $\left(1/D_{r},\ldots,1/D_{r}\right)$ this indicates that the particular component of the simplicial predictor does not carry too much statistical information. On the contrary, the further the coefficients associated with a component of the simplicial predictor are away from the barycentre, the more information this component carries. In this simple example, the third component does not seem very important, because the third row coefficients (B3) are close to the barycentre. The second component (B2) of the simplicial predictor on the other hand, is probably the most important component, because its vector of coefficients lie the furthest from the barycentre among the other three vectors of coefficients.
@@ 3.5 A test for specific values imposed on regression coefficients
A sub-composition stems from a composition by removing at least one component and rescaling the rest of the values to sum to unity. The resulting null hypothesis in this regression setting is $B_{j1}=\ldots=B_{jD_{r}}=0$ for some $j$, $1\leq j\leq D_{p}$. However, there is an associated problem with this, the rescaling part, which in order to be valid the property of sub-compositional coherence [^9] This property refers to the invariance of the results when the full composition or the sub-composition is used. (Aitchison, 2003) should be met and in this case it is not. However, the theoretical implications of the sub-compositionality are not known, and hence this test perhaps requires more investigation.
@@ 3.8.1 AR(1) model
Fiksel et al. (2022) mentioned that in the AR(1) model formulation, the current expected simplicial response is a linear function of its lagged time response, $E({\bf Y}_{t}|{\bf Y}_{t-1})={\bf Y}_{t-1}{\bf B}$, thus the matrix $\bf B$ can be interpreted as the matrix of transition probabilities of the states of the simplicial response at time $t-1$. The linear independence test in this case translates to testing the assumption of no auto-correlation. The case of allowing for more than one simplicial predictors extends the 1 lag to p lags, thus extend the AR(1) model to the AR(p) model. However, there is no theory to this end, what are implications of such a model (i.e. unit root testing, etc.) and thus this requires more research.
@@ 3.8.2 Categorical predictor or response
Similarly to the TFLR, the case of a categorical predictor can be addressed via the SCLS. In that case, each row of the simplicial predictor is written as ${\bf x}_{i}={\bf e}_{j}$, where ${\bf e}_{j}$ denotes a vector with 0 $D_{p}$ elements except its $j$-th element that is 1101010This could also be seen as an extreme case of compositional data, where one component takes a value of 1 and all other component have zero values., yielding a one-way analysis of variance type of model.
The linear independence test translates to testing the assumption of equal simplicial population means. The latter case was studied by Tsagris et al. (2017) using parametric and non-parametric procedures. However, simulation studies (not presented) replicating a case scenario from Tsagris et al. (2017) comparing the SCLS and TFLR with the bootstrap version of the two sample James test showed that the latter is size correct, but the former two produced inflated type I errors. Hence, in contrast to the suggestion made by Fiksel et al. (2022), there is evidence against the use of SCLS and TFLR models for the purpose of hypothesis testing of equality of simplicial means.
@@
3.9 Simplicial constrained median regression
@@
The simplicial response can be vectorized ${\bf y}^{*}=\mathop{\operator@font vec}\nolimits({\bf Y})$ and then estimate the coefficients of a logistic regression model of ${\bf y}^{*}$ on ${\bf X}^{*}$ with the identity as the link function, no constant term and imposing the same constraints as in the SCLS. This shows the equivalence between TFLR and univariate constrained logistic regression.
Further, this strategy is the same as the in one used in the seemingly unrelated regression equations (SURE) model (Zellner, 1962). The SURE idea can also be used if one wants to substitute the square loss (13), or the KLD loss (12) with the $L_{1}$ norm loss and thus estimate the regression coefficients by minimizing the $L_{1}$ norm between the observed and the fitted simplicial responses, $\sum_{i=1}^{n}\sum_{j=1}^{D_{r}}|y_{ik}-\hat{y}_{ij}|$, and thus perform what may be termed as simplicial constrained median regression. The median regression has been implemented in the quantreg R package (Koenker, 2023). That implementation allows for inequality constraints and hence the constraints will be nearly the same to those used in Eq. (14). Evidently, all aforementioned generalizations (adoption of the $\alpha$-transformation and inclusion of multiple simplicial predictors) are applicable in this case as well.
@@ 4.2 Type II error
The estimated powers for both models are presented in Table 2. For the small sample sized data, TFLR seems to produce higher estimated power levels than the SCLS, but as soon as the sample size increases their estimated power is equal  [^12] We also repeated the same experiment when the rows of the $\bf B$ matrix were nearly equal to $1/D_{r}$ but the power levels remained the same, indicating that the tests are really sensitive to small departures from the null hypothesis..